OpenStudy (anonymous):

multiply (5-3i)(6+4i)

OpenStudy (anonymous):

30+20i-18i-12i =30+2i-12i

OpenStudy (anonymous):

=30+20i-8i-12(i^2) =30+2i-12(i^2)